(1) The voltage of an electromagnetic wave traveling on a transmission line is given by V(z,t)=5e sin(41 X10t – 5772 +) a)
Answer #1

(1) V(2, 4) = 5e-43 in C42 X 10P / -5R2+7) a.) frequency =) f=8.5 = 42 X/07 (:w= 4a xiomad / ] 22 of = 2X107 If = 20942 wawecom 2) Amplitude of wave measured at I let amplitude be = XV. 26 = 5e-0.2554 X 10. x = se-2.554 R = 0.388 V d.) . Give Y (2)